Rank the following four acids in order of increasing bronsted acidity :
H2F+ , CH3OH, (CH3)2OH+ , CH3SH2+
Answer #1
Concepts and reason

Bronsted-Lowery acid is a proton donor and Bronsted-Lowery base is a proton acceptor. The strength of the acidity and basicity of the Bronsted-Lowery acids and bases depends on the polarity of bond between hydrogen and central atom.

First identify the central atom in the acidic species. Then compare the electronegativity difference between the hydrogen and the central atom, and compare the bond polarity of the species.

Fundamentals

According to Bronsted–Lowery theory, a chemical compound that donates H+{{\rm{H}}^ + } ion (proton) is known as Bronsted acid and a chemical compound that accepts H+{{\rm{H}}^ + } ion is called as Bronsted-Lowery base.

Central atom in the Bronsted acid H2F+{{\rm{H}}_2}{{\rm{F}}^ + } is fluorine, central atom in the acidic species CH3OH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{OH}} is oxygen, central atom present in the ion (CH3)2OH+{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{2}}}{\rm{O}}{{\rm{H}}^{\rm{ + }}} is oxygen and central atom of CH3SH2+{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{S}}{{\rm{H}}^{{\rm{2 + }}}} is sulfur.

In the Bronsted-Lowry acid H2F+{{\rm{H}}_2}{{\rm{F}}^ + } , fluorine is more electronegative than hydrogen. In the acidic species CH3OH{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{OH}} , (CH3)2OH+{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{2}}}{\rm{O}}{{\rm{H}}^{\rm{ + }}} , and CH3SH2+{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{S}}{{\rm{H}}^{{\rm{2 + }}}} the respective central atoms oxygen and sulfur are more electronegative than hydrogen atom. Fluorine is more electronegative than oxygen and sulfur.

The Bronsted acidity of the species is as follows:

H2F+>CH3SH2+>(CH3)2OH+>CH3OH{{\rm{H}}_2}{{\rm{F}}^ + }{\rm{ > C}}{{\rm{H}}_{\rm{3}}}{\rm{S}}{{\rm{H}}^{{\rm{2 + }}}}{\rm{ > }}{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{2}}}{\rm{O}}{{\rm{H}}^{\rm{ + }}}{\rm{ > C}}{{\rm{H}}_{\rm{3}}}{\rm{OH}}

Ans:

The increasing order of the Bronsted acidity is as follows:

H2F+>CH3SH2+>(CH3)2OH+>CH3OH{{\rm{H}}_2}{{\rm{F}}^ + }{\rm{ > C}}{{\rm{H}}_{\rm{3}}}{\rm{S}}{{\rm{H}}^{{\rm{2 + }}}}{\rm{ > }}{\left( {{\rm{C}}{{\rm{H}}_{\rm{3}}}} \right)_{\rm{2}}}{\rm{O}}{{\rm{H}}^{\rm{ + }}}{\rm{ > C}}{{\rm{H}}_{\rm{3}}}{\rm{OH}}